# Flow Snapshot Neurons in Action: Deep Neural Networks Generalize to Biological Motion Perception

### v1

@@ 4.4 Ablation Study of Hyper-parameters
To explore the effect of hyper-parameters in our model, ablation studies are conducted. Note that we use the same set of hyper-parameters for our model across all the datasets.
Number of slots K: As shown in Tab. 4, there is a non-monotonic performance trend versus the number of slots. The number of slots controls the diversity of temporal information captured from the path-level optical flow. An increase in the number of temporal slots can lead to redundant temporal clues, while a decrease might result in a lack of sufficient temporal clues.
The weight $\alpha$ of $L_{slot}$: The parameter $\alpha$ regulates the importance of $L_{slot}$ in comparison to other loss functions. A small value of $\alpha$ can impede the diversity in temporal slots, while a large $\alpha$ might adversely affect the optimization of features extracted from flow snapshot neurons and motion invariant neurons. From Tab. 4, it can be seen that when $\alpha$ is either small ($\alpha=1$) or large ($\alpha=100$), there will be a slight degradation in performance compared with our full model with $\alpha=10$.
Temperature $\tau$ in the patch-level optical flow: Tab. 4 demonstrates that higher temperature $\tau$ used for computing patch-level optical flows hurt performances, compared with our full model where $\tau=0.001$. This implies that temperature $\tau$ controls the smoothness of the sequence of flows and lower temperature $\tau$ is beneficial for clustering flows with similar movement patterns.
Temperature $\mu$ in $L_{slot}$: The temperature $\mu$ controls the sharpness of distribution over slots. High temperatures will negatively impact the convergence of slot feature extraction, while low temperatures can produce excessively sharp distributions, potentially impairing the stability of the optimization process during training. Tab. 4 indicates that configuring $\mu$ at either 0.02 or 0.20 are non-optimized options compared with our full model when $\mu$ is 0.05.
@@
Table 4: Ablation of key hyper-parameters in our model. Top-1 accuracy is reported on RGB videos, Joint videos with 6 points (J-6P), and SP videos with 8 points and a lifetime of 1 (SP-8P-1LT). From left to right, the ablated hyper-parameters are: the number of slots $K$ (Sec. 2.2)   the weight $\alpha$ of $L_{slot}$ (Sec. 2.2), temperature $\tau$ in the patch-level optical flow (Sec. 2.1) and temperature $\mu$ in $L_{slot}$ (Sec. 2.2). Best is in bold.
5 Discussion
Here, we introduce Motion Perceiver (MP) as a generalization model trained on natural RGB videos, capable of perceiving and identifying actions of living beings solely from their minimal motion patterns on point-light displays, even without prior training on such stimuli. Within MP, flow snapshot neurons learn prototypical flow patterns through competitive binding, while motion-invariant neurons ensure robustness to variations in temporal orders. The fused activations from both neural populations enable action recognition. To evaluate the generalization capabilities of all AI models, we curated a dataset comprising  63k stimuli across 24 BMP conditions using point-light displays inspired by neuroscience. Psychophysics experiments on this dataset were conducted, providing human behavioral data as an upper bound. By focusing solely on motion information, MP achieves superior generalization performance among all AI models and demonstrates alignment with human behavioral responses. Ethical concerns in video action recognition, such as privacy invasion, misuse of surveillance data, and potential algorithmic biases, can be alleviated by our method, which utilizes only patch-level optical flow information for action recognition. Despite promising generalization results in BMP stimuli, our model fails to consider attention and top-down influences. Furthermore, as it relies on patch-level optical flows alone as inputs, the integration of information from ventral (visual form perception) and dorsal (motion perception) pathways remains an open question.
Appendix A Detailed Results of All BMP Conditions
@@
The detailed results of all BMP conditions are shown in Tab. S1. It is evident that, in comparison to baselines, our motion perceiver (MP) not only delivers enhanced performance but also more accurately mirrors human performance in most BMP conditions. Especially in the case of SP-8P-1LT, our model significantly outperforms the top baseline by a substantial margin of 29.39%. These findings emphasize that our MP model is more effective at recognizing biological motion, demonstrating greater alignment with human behavioral data than existing AI models.
@@
Appendix B Viewpoint Discrepancy
As mentioned in Sec. 4.1, human behaviors in our study reveal a divergent trend compared to the human performance pattern described in [41]. We point out two reasons: first, the tasks differ between our study and [41]. Our study focuses on action recognition while the work [41] focuses on walking pattern discrimination. Second, we found the performance variations across different camera views depend on the action classes. As shown in Fig. S1, humans exhibit the best performance when viewing “Arm Circles" actions from the frontal view, because the movement of the two arms will not overlap in the frontal view. For the videos from the action class “Point To Something", the 90°view is the best viewpoint since it results in longer movement trajectories of the arm than other viewpoints. Therefore, although humans can achieve the best average accuracy on the 45°viewpoint in our experiment, it does not imply that humans always perform the best on the 45°viewpoint in all action classes.
Figure S1: Human performance across camera views within different action classes on J-6P videos in BMP dataset. The labels on the x-axis are the 10 action classes in the BMP dataset (Sec. 3.1).

### v2

@@ Appendix I Baselines with Optical Flow as Input
To investigate the impact of additional hyperparameters and key components in our model, we have conducted further ablation studies. Note that we use the same set of hyper-parameters for our model across all the datasets.
@@
Table S4: Ablation of key hyper-parameters in our model. Top-1 accuracy is reported on RGB videos, Joint videos with 6 points (J-6P), and SP videos with 8 points and a lifetime of 1 (SP-8P-1LT). From left to right, the ablated hyper-parameters are: the number of slots $K$ (Sec. 2.2)   the weight $\alpha$ of $L_{slot}$ (Sec. 2.2), temperature $\tau$ in the patch-level optical flow (Sec. 2.1) and temperature $\mu$ in $L_{slot}$ (Appendix, Sec. B.2). Best is in bold.
Temperature $\tau$ in the patch-level optical flow: Appendix, Tab. S4 demonstrates that higher temperature $\tau$ used for computing patch-level optical flows hurt performances, compared with our full model where $\tau=0.001$. This implies that temperature $\tau$ controls the smoothness of the sequence of flows and lower temperature $\tau$ is beneficial for clustering flows with similar movement patterns.
Number of slots K: As shown in Appendix, Tab. S4, there is a non-monotonic performance trend versus the number of slots. The number of slots controls the diversity of temporal information captured from the path-level optical flow. An increase in the number of temporal slots can lead to redundant temporal clues, while a decrease might result in a lack of sufficient temporal clues.
The weight $\alpha$ of $L_{slot}$: The parameter $\alpha$ regulates the importance of $L_{slot}$ in comparison to other loss functions. A small value of $\alpha$ can impede the diversity in temporal slots, while a large $\alpha$ might adversely affect the optimization of features extracted from flow snapshot neurons and motion invariant neurons. From Appendix, Tab. S4, it can be seen that when $\alpha$ is either small ($\alpha=1$) or large ($\alpha=100$), there will be a slight degradation in performance compared with our full model with $\alpha=10$.
Temperature $\mu$ in $L_{slot}$: The temperature $\mu$ controls the sharpness of distribution over slots. High temperatures will negatively impact the convergence of slot feature extraction, while low temperatures can produce excessively sharp distributions, potentially impairing the stability of the optimization process during training. Appendix, Tab. S4 indicates that configuring $\mu$ at either 0.02 or 0.20 are non-optimized options compared with our full model when $\mu$ is 0.05.
[…]
Table S5: Ablation of the time embedding (Sec. 2.4), the loss term $L_{slot}$ (Sec. 2.2) and the referenced frame for calculating path-level optical flow (Sec. 2.1) in our model. Top-1 accuracy is reported on RGB videos, Joint videos with 6 points (J-6P), and SP videos with 8 points and a lifetime of 1 (SP-8P-1LT). Best is in bold.
Model components: The ablation of the time embedding (Sec. 2.4) and the loss term $L_{slot}$ (Sec. 2.2) in our model is shown in Appendix, Tab. S5. As illustrated in Sec. 2.4, time embeddings are appended to the activations of flow snapshot neurons. Aligning with [66], time embeddings provide useful temporal context (A1). The loss $L_{slot}$ on contrastive walks encourages the slots to capture diverse prototypical flow patterns. Consistent with [102], removal of $L_{slot}$ (A2) leads to degraded generalisation performance, indicating the importance of contrastive walk loss between slots and optical flows. Furthermore, if our MP relies solely on the first frame as the reference for computing optical flow (A3), the performance will degrade significantly. This is because optical flows are estimated by computing the similarity between feature maps from video frames. The errors in feature similarity matching might be carried over in computing optical flows. Using multiple frames as references for computing optical flows eliminates such errors.
Downscale pixel-level flows: We also introduce an MP variation using pixel-level optical flow downscaled to the size of patch-level optical flow as input. Our MP model outperforms this model variation: 96.5% vs 68.8% in RGB, 69.0% vs 12.6% in J-6P, and 49.7% vs 9.4% in SP-8P-1LT. Hence, this implies that DINO captures semantic features that are more effective and robust for optical flow calculation than downscaled pixel-level flows.
Feature Extractor DINO with ResNet: In our MP, using ViT-based DINO has demonstrated its effectiveness in biological motion perception. We then conducted an experiment where ViT was replaced with the classical 2D convolutional neural network (2D-CNN) ResNet50, pre-trained on ImageNet, as a feature extractor for video frames. The results show that while the performance of our MP with ResNet50 is lower than with ViT, it still exceeds chance levels. Specifically, the accuracy of DINO-ViT (ours) versus DINO-ResNet50 is: 96.45% vs 80.37% in RGB, 69.00% vs 40.34% in J-6P, and 49.68% vs 40.03% in SP-8P-1LT. Additionally, it outperforms baselines using 3D-CNN backbones, such as ResNet3D, I3D, and R(2+1)D. This suggests that our MP effectively generalizes in BMP, regardless of the feature extraction backbone used.
Number of Trainable Parameters As mentioned in Sec. 4.3, we present the number of trainable parameters for all baselines and our MP model in Appendix, Tab. S2. While our model is larger than the baselines, its superior performance is not attributed to its size. To validate this, We included a SlowFast-ResNet101 variant with 61.9M parameters, which underperforms compared to our model. Specifically, the performance of our model versus SlowFast-ResNet101 is 96.5% vs 99.3% in RGB, 69.0% vs 39.4% in J-6P and 49.7% vs 12.6% in SP-8P-1LT. In addition, we list the number of trainable parameters in millions (M) for each model part of our MP: FlowSnapshot Neuron (0.07M), Motion Invariant Neuron (0M) and Feature Fusion (57.5M). Compared to DINO with 85.8M parameters for image processing, our MP model, appended to DINO, only requires slightly more than half of its size. Yet, it leverages DINO features from static images to generalize to recognize actions from a sequence of video frames.
Key Frame Analysis As indicated in Sec. 4.3, to explore how the start, end, or development of the action would influence action recognition, we analyze the effect of which frames are essential for the pick-up action class in one example video. Briefly, we randomly selected X frames among 32 frames, duplicated the remaining frames to replace these selected frames, and observed the accuracy drops, where X = [1,8,16,24,28,31]. When multiple frames are replaced, the performance drop implies the importance of the development of these frames. In total, we performed 1000 times of random frame selections per X and presented the visualization of frame importance by averaging all the accuracy drops over all the random frame selections. The visualization results in Appendix, Fig. S6 suggest that the fourth and seventh frames are essential for the pick-up class recognition.
Figure S6: Frame analysis for the "pick-up" action class in one example video from the BMP dataset. The color in the heat map indicates the video frames which would impact the action recognition accuracy of the model the most. Blue colors represent higher importance.
Appendix L Detailed Results of All BMP Conditions
@@
The detailed results of all BMP conditions are shown in Appendix, Tab. S6. It is evident that, in comparison to baselines, our motion perceiver (MP) not only delivers enhanced performance but also more accurately mirrors human performance in most BMP conditions. Especially in the case of SP-8P-1LT, our model significantly outperforms the top baseline by a substantial margin of 29.39%. These findings emphasize that our MP model is more effective at recognizing biological motion, demonstrating greater alignment with human behavioral data than existing AI models.
@@
Appendix M Social Impact
The Motion Perceiver, with its human-like ability to perceive biological motion, brings both promising advancements and concerning implications for society. In the sports and fitness field, it provides valuable tools for improving athletic performance, injury prevention, and personalized training. However, the same capabilities may raise privacy concerns, as the technology could potentially be misused for unauthorized surveillance or tracking of individuals’ movements in public spaces. This invasion of privacy might extend to analyzing behavior patterns and gait. Furthermore, there is a risk of discrimination if the method is employed in contexts like employment screening or law enforcement profiling, where certain movement patterns might be unfairly associated with specific demographics or lead to biased judgments.